OpenStudy (anonymous):

?? 3 over x plus 13 over quantity x minus 11 & 5 over 4 x squared y minus y over 14 x z

OpenStudy (anonymous):

1) \[\frac{ 3 }{ x }+\frac{ 13 }{ x-11 }\]

OpenStudy (anonymous):

2) \[\frac{ 5 }{ 4x^2y }-\frac{ y }{ 14xz }\]

OpenStudy (nurali):

\[\frac{ 3 }{ x }+\frac{ 13 }{ x-11 }\] \[\frac{ 3(x-11)+13(x) }{ (x)(x-11) }\] \[\frac{ 3x-33+13x }{ x^2-11x }\] \[\frac{ 16x-33 }{ x^2-11x }\]

OpenStudy (nurali):

\[\frac{ 5 }{ 4x^2y }-\frac{ y }{ 14xz }\] \[\frac{ 5(14xz)-y(4x^2y) }{ (4x^2y)(14xz) }\] \[\frac{ 70xz-4x^2y^2 }{ 56x^3yz }\] \[\frac{ 70z-4xy^2 }{ 56x^2yz }\]